OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 8 feet and 6 feet and the height is 10 feet?

OpenStudy (anonymous):

use 2(wh+lw+lh)

OpenStudy (anonymous):

how do I plug that in?

OpenStudy (anonymous):

w=width l=length h=height

OpenStudy (anonymous):

so 480

OpenStudy (anonymous):

divided by 2?

OpenStudy (anonymous):

which is 240

OpenStudy (anonymous):

@Hero @SmoothMath @timo86m ??

OpenStudy (anonymous):

could you hellp?

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

i'm just wondering if 240 is right

OpenStudy (anonymous):

thanks

hero (hero):

LA = Perimeter of Base x Height

OpenStudy (anonymous):

so then what is it?

hero (hero):

LA = 2(l + w)h

hero (hero):

LA = 2(l + w)h LA = 2(8 + 6)10 LA = 2(14)10 LA = 28∙10 LA = 280

OpenStudy (anonymous):

thank you soso much!

OpenStudy (anonymous):

What is the lateral area of a triangular prism with base edges all equaling 3 meters and a height of 5 meters?

OpenStudy (anonymous):

is the answer to that 45

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What is the area of each lateral face?

OpenStudy (anonymous):

could you explain

jimthompson5910 (jim_thompson5910):

Here's a drawing that might help 

jimthompson5910 (jim_thompson5910):

The lateral faces are the ones I'm going to shade below